Assume a Company Xs stock price today and cost of equity is consistent with the dividend discount model for constant growth; HINT: Cost of Equity = Dividend Yield + Capital Gain Yield. Use the information in the table below to answer problems 16-17. Any missing information will need to be solved for. TIME (Years) 0 N 1 350.00 Dividends Shares Outstanding Book Value Price Per Share Return On Equity Cost of Equity Long Term Growth Dividends Per Share 135.00 $6,750.00 80.00 12.0% 16. Company Xs dividend payout ratio is closest to: A. 0.3 B. 0.4 C. 0.5 D. 0.6 E. 0.7 17. The present value of Company Xs first dividend payment one year from now (T1) is closest to: A. $3.33 B. $3.40 C. $3.73 D. $3.81 E. $3.83

Questions 16 & 17
Answer #1
Return on equity = 1350/6750 = 20.00%
Growth = ROE*Retention
0.12 = 0.20*Retention
Retention = 0.12/0.20 = 60%
Dividend payout = 1-Retention ratio = 1-60% = 40.00%
Dividends = 1350*40% = $           540.00
Dividend per share = 540/135 = $                4.00
Cost of equity = 4*1.12/80+0.12 = 17.60%
16) Dividend payout = 40.00% [Option B]
17) PV = 4*1.12/1.176 =$3.81 [Option D]